2x2-5x+7=3 by completing the square
Answer by ewatrrr(10682)


Hi,
2x2-5x+7=3 by completing the square
2x2-5x+4=0
x^2 -(5/2)x + 2
x^2 -(5/2)x + 25/16 -25/16 + 2 = 0
[x^2 -(5/2)x + 25/16] -25/16 + 32/16 = 0
(x - 5/4)^2 + 7/16 = 0